# Perception of Visual Content: Differences between Humans and Foundation Models

### v2

@@ Conclusion
Answering the RQs. We first look at RQ1: “How similar are human-generated and ML-generated annotations?”. It can be concluded that from a low-level perspective, i.e., what types of word occur, sentence structures, the ML Captions and Human Labels are similar to each other, but ML Objects are relatively different to either of them. This is expected due to similarities in underlying cognitive processes between ML models and human brains. However, when looking at them from a high-level point of view, despite the apparent low similarity between ML Objects and the other annotations, all of them are alike in that they exhibit consistent similarity or dissimilarity, even when considering the spread based on region, where all of them have the same trends, with Europe clustering together and showing a strong agreement across all annotation types. This led to RQ2: “How do different combinations of annotations affect fairness in ML predictive models?” In the case of region classification, it was revealed that different annotation combinations work best for different geographical regions and image categories. This points to the conclusions that all annotations are important, and machine-generated annotations cannot just replace human-generated ones. It is additionally apparent that each region obtained best classification performance with models that utilised annotation combinations that included ML Captions. In the case of income regression, the combination of ML Objects and ML Captions is what results in enhanced performance. Unlike the region classification, the best performance for individual classes was obtained from different annotation models. Considering these findings, it can be concluded that there is no particular combination of annotations that result in best region-based performance as that is highly dependent on the quality of the images and how well the human annotator or machine is able to select the best discriminative features.
@@
Implications. These findings have several implications. First, despite their apparent differences, humans and machines may be more alike in their bias, or lack thereof, in the context of visual content. Their consistent similarity and dissimilarity of annotations across regions implies that their level of bias is comparable. This insight could steer researchers towards further advancing ML models to avoid as much bias as possible.
@@
Second, due to machine annotations (ML Captions only and ML Objects-ML Captions) performing better overall at both the classification and regression tasks, tasks in other domains could potentially benefit from using machine annotations as well. However, this could result in a potential negative societal impact in that humans become overly reliant on machines, which may cause them to miss possible biases in the models that originated from the training data, miss errors due to assuming the superior performance of the ML models, or even cause difficulties in identifying accountability if the ML models’ prediction errors result in significant repercussions. This could also lead to potential misuse of the technology as it is possible that one intentionally uses a model that is potentially biased without attempting to address the underlying issues.

### v3

@@ Conclusion
Answering the RQs. We first look at RQ1: “How similar are human-generated and ML-generated annotations?”. From a low-level perspective, i.e., what types of word occur and sentence structures, the ML Captions and Human Labels are similar to each other, but ML Objects are relatively different to either of them. This is expected due to similarities in underlying cognitive processes between ML models and human brains. However, from a high-level point of view, despite the apparent low similarity between ML Objects and the other annotations, all exhibit consistent trends, even when considering the spread based on region, with Europe clustering together and showing a strong agreement across all annotation types. This led to RQ2: “How do different combinations of annotations affect the performance and bias in ML predictive models?” For region classification, different annotation combinations work best for various regions and image categories. Models that included ML Captions achieved the best performance in each region. Most image categories performed best with annotations containing at least two types. For income regression, using ML Object Labels and ML Captions enhanced performance, particularly across many image categories. The best performance for each region was obtained from a wide variety of annotations. Similar to region classification, performance was highest in Africa when all annotation types were used, likely due to the images containing object/action cues and nuanced environmental context. ML-based annotations worked well for action image categories, while human input benefited non-action ones. Using diverse annotations led to more consistent predictions, as seen in Africa’s lower variance in the scatter plot. Considering these findings, it can be concluded that all annotations are important and machine-generated annotations cannot fully replace human-generated ones.
@@
Implications. These findings have several implications. First, despite their apparent differences, humans and machines may share similar levels of bias in visual content. Their consistent patterns across regions show comparable tendencies. This points to the need for further advancement to reduce bias in ML models.
@@
Second, since machine annotations (ML Captions only and ML Objects-ML Captions) perform better overall at both the classification and regression tasks respectively, they may be beneficial in other domains. However, being overly reliant on machine annotations may have negative societal impacts, such as missing possible biases in the models that originated from the training data, assuming ML models are always superior, or struggling to identify accountability if errors lead to significant consequences. This could also lead to potential misuse of the technology, where biased models are intentionally used without addressing underlying issues.